OpenStudy (anonymous):

Find the surface area of a cylinder in terms of pi. The diameter is 14 and the height is 18

OpenStudy (anonymous):

Curved surface area or the total surface area?

OpenStudy (anonymous):

Total

OpenStudy (anonymous):

Alright! Now, first the formula : \[2\pi r ( r + h)\] r = 14/2 = 7 h = 18 Substitute all the values, and you'll get : \[2\times \frac{ 22 }{ 7 } \times 7 ( 7 + 18) \]

OpenStudy (anonymous):

You should be able to solve that, right?

OpenStudy (anonymous):

Yes thank you very much